Moor Company uses a job-order costing system and applies overhead costs
to jobs at a rate of 120% of direct labor cost. Moor Company started the
month of June with three jobs in process. The cost of beginning work in
process plus the costs added during June are shown below:

                                  Job #1-A      Job #3-G      Job #7-L
beginning work in process .....    $3,250        $2,690        $3,720

COSTS ADDED DURING JUNE:
direct materials ..............    $4,680        $3,950        $2,870
direct labor ..................     3,530         4,180         2,410

During June, Moor Company completed and sold Job #1-A. Job #3-G was
also completed but was not sold by the end of June. Job #7-L was not
completed by the end of June. Moor Company's cost of goods sold for
June after the overhead variance was closed totaled $16,290.

Calculate Moor Company's actual overhead cost for the month of June.
Answer #1
Job #1-A
Beginning work in process 3250
Direct materials 4680
Direct labor 3530
Overhead applied 4236 =3530*120%
Total cost of Job #1-A 15696
Total Direct labor during June 10120 =3530+4180+2410
X Overhead rate 120%
Overhead applied 12144
Overhead applied 12144
Add: Underapplied overhead 594 =16290-15696
Actual overhead cost for the month of June 12738